# Adaptive Energy Regularization for Autonomous Gait Transition and Energy-Efficient Quadruped Locomotion

### v1

@@ Abstract
In reinforcement learning for legged robot locomotion, crafting effective reward strategies is crucial. Pre-defined gait patterns and complex reward systems are widely used to stabilize policy training. Drawing from the natural locomotion behaviors of humans and animals, which adapt their gaits to minimize energy consumption, we propose a simplified, energy-centric reward strategy to foster the development of energy-efficient locomotion across various speeds in quadruped robots. By implementing an adaptive energy reward function and adjusting the weights based on velocity, we demonstrate that our approach enables ANYmal-C and Unitree Go1 robots to autonomously select appropriate gaits—such as four-beat walking at lower speeds and trotting at higher speeds—resulting in improved energy efficiency and stable velocity tracking compared to previous methods using complex reward designs and prior gait knowledge. The effectiveness of our policy is validated through simulations in the IsaacGym simulation environment and on real robots, demonstrating its potential to facilitate stable and adaptive locomotion. Videos and more details are at https://sites.google.com/berkeley.edu/efficient-locomotion
@@ I Introduction
Humans and animals exhibit various locomotion behaviors at different speeds, optimizing for their energy efficiency. For instance, humans typically walk at low speeds and run at higher speeds, rarely opting for jumping. Prior research demonstrated through optimal control on planar models the correlation between speed and optimal gait choices concerning the cost of transport (CoT). For quadrupeds, the optimal gaits were four-beat walking [^1] Four-beat walking, two-beat walking, trotting and galloping are typical gaits for quadruped robots defined in [1] based on feet contact schedule. at low speeds, trotting at intermediate speeds, and trotting/galloping at high speeds [1].
@@
Due to the rich information in the gaits, using a gait as guidance for locomotion policies is popular among lots of reinforcement learning (RL) based methods [2, 3]. However, crafting a versatile and robust locomotion policy that can adapt to and transition between multiple speeds while generalizing across different platforms poses substantial challenges. One of the main challenges here is the reward design. Gait reference can be used as extended state or extra regularization terms in reward functions to provide more supervision. Previous works [3, 4, 5] trained on different quadruped robots within simulation environments like IsaacGym [6, 7] and successfully transferred these policies to physical hardware. However, they often necessitate intricate reward designs and weight tuning. Apart from gait information, reward terms like feet-air time and contact force penalizing [4] were also used to encourage specific behaviors and help stabilize the training. While these additional reward components are aimed at inducing or preventing specific behavioral traits, they inadvertently align with the broader objective of reducing energy costs. This convergence prompts a reconsideration of our reward strategy: could a more straightforward, energy-centric reward term be used to replace the specifically designed terms used in prior policy training? Such a term would encapsulate the core objective of reducing energy consumption and fostering stable and efficient locomotion.
@@
Figure 1: Compared to legged-gym baseline [4], our single policy (from one-time RL training) autonomously adopted different energy-efficient gaits (four-beat walking and trotting). It achieved lower energy consumption (low leg-swing) at varying speeds ($0.1m/s$ interval; mean and variance from 50 independent runs).
@@
Building on the concept that energy-efficient gaits correlate with speed [1] and aligning with prior work emphasizing that energy minimization at pre-selected speed results in the emergence of specific gaits [8], this study investigates a more streamlined reward formulation for energy-efficient locomotion. By focusing on energy minimization without intricately designed reward components, we aim to verify if such a simplified approach can yield stable and effective velocity-tracking in quadruped robots across various speeds. Instead of generating multiple velocity-specific energy optimal policies [8], we focus on getting a single energy-optimal policy across all target velocities via RL training.
In our research, we examine the influence of energy regularization weights on policy performance, identifying that both excessively low and high weights can lead to unnatural movements or immobility. Recognizing that energy terms have different scales across velocities and require adaptive velocity-conditioned weights, we first design a non-negative energy reward function and then find an adaptive reward form by interpolating the maximum energy weights at selected speeds to facilitate effective velocity tracking.
@@
Employing this adaptive reward structure within IsaacGym enables the training of robust policies for the ANYmal-C [9] and Unitree Go1 [10] quadruped robots. Our methodology, illustrated in Figure 1, identifies appropriate gaits, such as four-beat walking at lower speeds and trotting at higher speeds, without predefined gait knowledge—outperforming baseline approaches [4] in energy efficiency. Our policy also significantly improves velocity tracking and energy consumption performance compared to a policy trained with fixed-weight energy rewards. The trained single policy is deployed on a real Go1 robot to verify its stable moving and transition locomotion skills in the real world.
@@
  • Introduction of a streamlined reward formula integrating basic velocity-tracking and adaptive energy minimization to foster stable, velocity-sensitive locomotion policies.
@@
  • Evaluation of the velocity-tracking and energy efficiency across reward structures and weight tunings for ANYmal-C and Unitree Go1, culminating in the real-world application of these policies on a Go1 robot, affirming their efficacy in stable locomotion and gait transition.
@@ II-A Reinforcement Learning for Locomotion Skills
Many efforts were made to overcome this limitation by promoting the behavioral diversity of legged robots. Researchers in [20] proposed a hierarchical framework that pre-specifies a set of gait primitives and allows an RL model to choose from them. In [21, 22], gait primitives were parameterized using the contact schedule and RL policy was trained to select these parameters. These works used model predictive control (MPC) as the lower-level controller, which demands preliminary knowledge of locomotion and contact modeling [23]. In [3], behavioral-related arguments such as body height, step frequency, and phase are directly added to the RL model input for end-to-end training. Although various two-beat gaits (where legs touch the ground in pairs) were realized, they cannot generate four-beat gaits (where legs touch the ground in orders) and require manual gait specification in different scenarios. It is more desirable if the quadruped robot can select the most suitable behavior by itself. In [8], a pipeline was proposed to output different gaits under different velocities via minimizing energy consumption. However, this pipeline relies on training and distillation of several velocity-specific RL policies.
@@ II-B Energy Studies on Locomotion
Figure 2: Top row: Gait plot generated from the ANYmal-C simulation where the legged robot is first commanded to move at $0.5$ m/s from 00 to $2$ seconds and demonstrated four-beat walking (four legs touch the ground one by one in order, refer to [1] for details), then at $1.0$ m/s from $2$ to $4$ seconds and demonstrated a transition gait between walking and trotting, finally at $2.0$ m/s for $4$ to $6$ seconds and demonstrated trotting (legs touch the ground in pairs). Middle row: Snapshots from the ANYmal-C simulation taken approximately at the vertical purple lines in the gait plot. Bottom row: Snapshots of Go1 moving on a playground. Its feet contact state approximately corresponds to the vertical purple line in the gait plot.
III Locomotion Reward Design
@@
$R=R_{motion}+R_{energy}+R_{others}$ (1)
where $R_{motion}$ encourages accurate velocity tracking, $R_{energy}$ discourages energy consumption and $R_{others}$ includes other necessary rewards to stabilize training. In previous work [8], motion rewards include penalty on linear and angular velocity tracking errors; energy rewards include penalty on motor power with a fixed weight; survival bonus is also added. In experiments, we found the training process unstable potentially due to the negative nature of tracking and energy rewards. Besides, each energy reward weight usually only works within a very narrow range of reference speeds. It is hard to find a single energy reward weight value that works for all reference velocities without knowing more simulation settings and training details. As a result, we proposed the following reward function to promote the automatic generation of energy-efficient behavior of legged robots under various reference velocities.
$R=\frac{1}{Z(\hat{v}_{x})}\Big{[}R_{lin}+\alpha_{ang}R_{ang}+\alpha(\hat{v}_{x})R_{en}\Big{]}$ (2)
where $\alpha_{ang}=0.5$, $\hat{v}_{x}$ is the user-specified reference velocity, $\alpha(\hat{v}_{x})$ is the adaptive energy reward weight, $Z(\hat{v}_{x})$ is the normalizing index for total reward, $R_{lin},R_{ang}$ are velocity tracking rewards, and $R_{en}$ is the energy reward. The remaining section elaborates on each component in (2).
@@ III-1 Motion Rewards
$R_{lin}$ and $R_{ang}$ respectively encourage the legged robot to track the linear reference velocities in two directions $\hat{v}_{x},\hat{v}_{y}$ and angular reference velocities $\hat{\omega}_{z}$.
$\begin{split}R_{lin}&=\exp\Big{(}-\frac{|v_{x}-\hat{v}_{x}|^{2}+|v_{y}-\hat{v}_{y}|^{2}}{\sigma_{v}}\Big{)}\\ R_{ang}&=\exp\Big{(}-\frac{|\omega_{z}-\hat{\omega}_{z}|^{2}}{\sigma_{\omega}}\Big{)}\end{split}$ (3)
$\hat{v}_{y}$ and $\hat{\omega}_{z}$ are not user-specified commands, but randomly sampled during training as explained in section IV. $\sigma_{v}$ and $\sigma_{\omega}$ are scaling factors depending on the training velocity range. The structure of motion rewards, the coefficient $\alpha_{ang}=0.5$ for angular velocity tracking, and the scaling coefficients follow the default setting in legged-gym [4].
@@ III-2 Energy Rewards
$R_{en}=\exp\Big{(}-\frac{\sum_{i}|\tau_{i}||\dot{q}_{i}|}{\sigma_{en}}\Big{)}$ (4)
The exponential form guarantees a positive reward. $\tau$ are each joint’s actuated torques, and $\dot{q}$ are joint velocities. We multiply the absolute values of each entry of $\tau$ with each entry of $\dot{q}$ and sum them up in (4) to follow the fact that a motor does not get charged back even when the applied torque is opposite to the motion [32]. $\sigma_{en}$ is an energy scaling constant.
III-3 Adaptive Energy Weight
$\alpha(\hat{v}_{x})$ is the weight of the energy reward terms. Previous works take this value as a constant [8, 32], but we argue that it should be adaptive to the reference velocity $\hat{v}_{x}$ to achieve suitable behaviors directly from RL. To verify this, we first run RL under fixed sampled $\hat{v}_{x}$ using pre-selected $\sigma_{en}$ and try various values of $\alpha(\hat{v}_{x})$. When $\alpha(\hat{v}_{x})$ is too large, the robot tends to stay unmoved to save energy, neglecting the velocity tracking task. When $\alpha(\hat{v}_{x})$ is too small, the robot tends to use a highly inefficient and unnatural way to walk.
To find an adaptive weight, we first collect the largest $\alpha(\hat{v}_{x})$ when the velocity tracking error is smaller than a pre-defined small threshold $\delta$ for a set of speeds $\hat{v}_{x}$. After collecting velocity-weight sample pairs $(\hat{v}_{x,j},\alpha(\hat{v}_{x,j}))_{j=1}^{M}$ for a range of velocities, we see a clear trend (see Figure 3): with velocity increasing, the maximum allowed $\alpha$ decreases. This trend corresponds to the fact that the kinetic energy increases quadratically with the velocity. While the motion rewards are expected to converge close to zero, the energy reward has variant optimal values across velocities. Therefore, we can set a larger weight for lower-speed training but need to decrease it as the velocity increases. When training velocity-conditioned locomotion policies, we linearly interpolate between selected pairs to acquire the velocity-conditioned energy weight $\alpha(\hat{v}_{x})$ for the current sampled velocity. We will show in experiments that this is a simple but effective way of training energy-effective policies across different velocities.
III-4 Normalization Index
$Z(\hat{v}_{x})$ is an adaptive normalization index. Due to the adaptive $\alpha(\hat{v}_{x})$, the reward scale under different command velocities varies. This often leads to instability in RL training. As such, for each velocity-weight sample pair $(\hat{v}_{x,j},\alpha(\hat{v}_{x,j}))$, we also record the final achieved reward $Z(\hat{v}_{x,j})$ and analogously use linear interpolation to get the normalization index curve $Z(\hat{v}_{x})$ to stabilize training.
IV Locomotion Skill Training Details
Section III focused on the reward design of the proposed method, which constitutes the most essential part of RL training. However, the energy reward is not a stand-alone one that can be used directly for RL training, and we have to use it together with the basic locomotion rewards. Across quadruped locomotion baselines, the default training settings and locomotion rewards slightly differ. This section explains the basic RL training and simulation settings other than energy regularization. These basic settings can also generate a naive locomotion policy without energy regularization, but these policies usually have low energy efficiency and might have undeployable abnormal behavior. The two baselines we use in this research are legged-gym [4] for ANYmal-C and walk-these-ways [3] for Unitree Go1.
IV-A ANYmal-C Settings

### v2

@@ Abstract
In reinforcement learning for legged robot locomotion, crafting effective reward strategies is crucial. Predefined gait patterns and complex reward systems are widely used to stabilize policy training. Drawing from the natural locomotion behaviors of humans and animals, which adapt their gaits to minimize energy consumption, we investigate the impact of incorporating an energy-efficient reward term that prioritizes distance-averaged energy consumption into the reinforcement learning framework. Our findings demonstrate that this simple addition enables quadruped robots to autonomously select appropriate gaits—such as four-beat walking at lower speeds and trotting at higher speeds—without the need for explicit gait regularizations. Furthermore, we provide a guideline for tuning the weight of this energy-efficient reward, facilitating its application in real-world scenarios. The effectiveness of our approach is validated through simulations and on a real Unitree Go1 robot. This research highlights the potential of energy-centric reward functions to simplify and enhance the learning of adaptive and efficient locomotion in quadruped robots. Videos and more details are at https://sites.google.com/berkeley.edu/efficient-locomotion
@@ I Introduction
Humans and animals exhibit various locomotion behaviors at different speeds, optimizing for their energy efficiency. For instance, humans typically walk at low speeds and run at higher speeds, rarely opting for jumping [1]. Prior research demonstrated through optimal control on planar models the correlation between speed and optimal gait choices concerning the cost of transport (CoT). For quadrupeds, the optimal gaits were four-beat walking [^1] Four-beat walking, two-beat walking, trotting and galloping are typical gaits for quadruped robots defined in [2] based on feet contact schedule. at low speeds, trotting at intermediate speeds, and trotting/galloping at high speeds [2].
@@
Due to the rich information in the gaits, using a gait as guidance for locomotion policies is popular among lots of reinforcement learning (RL) based methods [3, 4]. However, crafting a versatile and robust locomotion policy that can adapt to and transition between multiple speeds while generalizing across different platforms poses substantial challenges. One of the main challenges here is the reward design. Gait reference can be used as extended state and extra regularization terms in reward functions to provide more supervision or input for low-level MPC [5] controllers. Previous works [4, 6, 7] trained on different quadruped robots within simulation environments like IsaacGym [8, 9] and successfully transferred these policies to physical hardware. However, they often necessitate intricate reward designs and weight tuning. Apart from gait information, reward terms like feet-air time and contact force penalizing [6] were also used to encourage specific behaviors and help stabilize the training. While these additional reward components are aimed at inducing or preventing specific behavioral traits, they inadvertently align with the broader objective of reducing energy costs [10, 11]. This convergence prompts a reconsideration of our reward strategy: Could a more straightforward, energy-centric reward term replace the specifically designed terms used in prior policy training and work for general locomotion tasks, including various linear and angular velocity tracking and terrains? Such a term would encapsulate the core objective of reducing energy consumption and fostering stable and efficient locomotion. Gaits at different speeds and terrains are generated posteriorly when the robot finds an energy-efficient way to solve the locomotion task.
@@
Figure 1: Compared to the baseline when there is no energy regularization, our single policy (from one-time RL training) autonomously adopted different energy-efficient gaits (walking, trotting and fly trotting). It achieved lower energy consumption (adjusted leg-swing) at varying speeds.
Figure 2: Gait switching under different command velocities. The policy is generated when $\alpha_{en}=1.0$. As the command velocity increases, the policy shows automatic gait transition. We also demonstrate snapshots of two beat walking at 0.5 m/s, trotting at 1.4 m/s and fly trotting at 2.3 m/s.
@@
Building on the concept that energy-efficient gaits correlate with speed [2] and aligning with prior work emphasizing that energy minimization at pre-selected speed results in the emergence of specific gaits [11], this study investigates a more streamlined reward formulation for energy-efficient locomotion. By focusing on energy minimization without intricately designed reward components, we aim to verify if such a simplified approach can yield stable and effective velocity-tracking in quadruped robots across various speeds. Instead of generating multiple velocity-specific energy optimal policies [11], we focus on getting a single energy-optimal policy across all target linear and angular velocities, and different terrains via RL training.
In our research, we investigate the impact of incorporating a distance-averaged energy reward term into the reinforcement learning framework. This reward term directly penalizes energy consumption per unit motion traveled, promoting energy-efficient locomotion across various speeds. We explore the effect of different weightings for this energy reward, observing that both excessively low and high weights can lead to undesirable behaviors such as unnatural movements or immobility. By carefully tuning the weight of the distance-averaged energy reward, we demonstrate its effectiveness in facilitating stable velocity tracking and encouraging the emergence of energy-efficient gaits.
@@
Employing this adaptive reward structure within IsaacGym enables the training of robust policies for the Unitree Go1 [12] quadruped robot. As illustrated in section IV, our methodology identifies appropriate gaits, such as four-beat walking at lower speeds and trotting at higher speeds, without predefined gait knowledge. The energy-efficient policy also shows better performance in circle tracking and terrain clearance tasks. The trained single policy is deployed on a real Go1 robot to verify its stable moving and transition locomotion skills in the real world.
@@
  • Introduction of a streamlined reward formula integrating basic velocity-tracking and distance-averaged energy minimization to foster stable, velocity-sensitive locomotion policies.
@@
  • Evaluation of the velocity-tracking and energy efficiency across reward structures and weight tunings for Unitree Go1, culminating in the real-world application of these policies on a Go1 robot, affirming their efficacy in stable locomotion and gait transition.
@@ II-A Reinforcement Learning for Locomotion Skills
Many efforts were made to overcome this limitation by promoting the behavioral diversity of legged robots. Researchers in [22] proposed a hierarchical framework that pre-specifies a set of gait primitives and allows an RL model to choose from them. In [5, 23], gait primitives were parameterized using the contact schedule and an RL policy was trained to select these parameters. These works used model predictive control (MPC) as the lower-level controller, which demands preliminary knowledge of locomotion and contact modeling [24]. In [4], behavioral-related arguments such as body height, step frequency, and phase are directly added to the RL model input for end-to-end training. Although various two-beat gaits (where legs touch the ground in pairs) were realized, they cannot generate four-beat gaits (where legs touch the ground in orders) and require manual gait specification in different scenarios. It is more desirable if the quadruped robot can select the most suitable behavior by itself. In [11], a pipeline was proposed to output different gaits under different velocities via minimizing energy consumption. However, this pipeline relies on training and distillation of several velocity-specific RL policies.
@@ II-B Energy Studies on Locomotion
III Energy Regularization
@@
$R=(R_{motion}+R_{energy})*f(R_{aux}),$ (1)
where $R_{motion}$ encourages accurate velocity tracking, $R_{energy}$ discourages energy consumption and $R_{aux}$ includes other necessary rewards to stabilize training. * and $f$ are respectively an arithmetic operator and functions. Common choices are =+*=+∗ = +, $f(x)=-x$ [6] and =×*$\times$ = ×, $f(x)=\exp(-x)$ [4].
In previous work [11], motion rewards include negative squared linear and angular velocity tracking errors; energy rewards include negative time-averaged motor power with a fixed weight $-\tau\dot{q}$; survival bonus is also added. In experiments, we found such a training process unstable across different speeds, primarily for two reasons—1) The negatively unbounded nature of tracking and energy rewards. 2) The scale of energy reward varies across different speeds, and the energy reward weight usually only works within a narrow range of reference speeds. It is hard to find a single energy reward weight value that works for all reference velocities without knowing more simulation settings and training details. As a result, different speeds are trained in different runs separately in [11] to energy-efficient gaits at different speeds.
(a) Straight Line Walking
(b) Angular Spining
(c) Terrain Walking
(d) Anymal-C Walking
Figure 3: Ablation study of energy consumption in Unitree Go1 simulation. For straight line walking, reference linear velocities are chosen from $0.1$ to $2.5$ m/s with $0.1$ common gap. The cost of transportation is measured in $J/m$. For angular spining, reference angular velocities are chosen from $-2.5$ to $2.5$ rad/s with $0.2$ common gap. In both (a) and (b), CoT considerably decreases when $\alpha_{en}$ reaches $1.0$. CoT of $\alpha_{en}=1.5$ when reference velocity of above $1.9$ m/s is not plotted because the output velocity drops to zero in this range. This indicates that velocity tracking accuracy will be sacrificed when energy regularization weight $\alpha_{en}$ is too large. For terrain walking, the robot is asked to walk in a straight line on a rough slope terrain with reference linear velocities from $0.3$ to $1.5$ m/s with $0.1$ common gap, because it is hard to walk either too slowly or too quickly on such terrains. (c) shows that the reduced CoT analogously appears on terrains. (d) shows the effect of energy regularization method to ANYmal-C platform with the same parameters $\sigma_{en,x}$, $\sigma_{en,z}$ and $\alpha_{en}$.
To overcome this challenge in reward design, we proposed that the energy-related reward function should depend on the robot’s velocity to promote an automatic generation of energy-efficient behavior of legged robots under various reference velocities.
$R=(R_{motion}+\alpha_{en}R_{en}(v_{x},\omega_{z}))\exp(-R_{aux})$ (2)
where $v_{x}$ is the robot moving velocity. The remaining of this section elaborates on each component in (2).
@@ III-1 Motion Rewards
$R_{motion}$ consists of $R_{lin}$ and $R_{ang}$, which respectively encourage the legged robot to track the linear reference velocities in two directions $\hat{v}_{x},\hat{v}_{y}$ and angular reference velocities $\hat{\omega}_{z}$.
$\begin{split}&R_{motion}=R_{lin}+\alpha_{ang}R_{ang},\\ &R_{lin}=\exp\Big{(}-\frac{|v_{x}-\hat{v}_{x}|^{2}+|v_{y}-\hat{v}_{y}|^{2}}{\sigma_{v}}\Big{)},\\ &R_{ang}=\exp\Big{(}-\frac{|\omega_{z}-\hat{\omega}_{z}|^{2}}{\sigma_{\omega}}\Big{)},\end{split}$ (3)
where $\hat{v}_{y}$ and $\hat{\omega}_{z}$ are not user-specified commands, but randomly sampled during training as explained in section IV-A. $\sigma_{v}$ and $\sigma_{\omega}$ are scaling factors depending on the training velocity range. The structure of motion rewards, the coefficient $\alpha_{ang}=0.5$ for angular velocity tracking, and the scaling coefficients follow the default setting in legged-gym [6].
@@ III-2 Energy Rewards
$R_{en}=\exp\Big{(}-\frac{\sum_{i}|\tau_{i}||\dot{q}_{i}|}{\sigma_{en,x}|v_{x}|+\sigma_{en,z}|\omega_{z}|}\Big{)}$ (4)
The energy consumption is averaged against the robot’s amount of motion. $\tau_{i}$’s are the actuated torque at each joint while $\dot{q}_{i}$’s denotes the joint velocities. $\sigma_{en,x}$ and $\sigma_{en,z}$ are energy scaling constants. Components inside the round bracket of (4) is equivalent to the linear-distance-averaged energy consumption when $\sigma_{en,z}=0$, which is the conventional definition of the cost of transportation (CoT). Since we are training a linear and angular velocity-dependent policy, we include rotation distance to form a generalized distance while generating the energy reward.
As discussed in [5], directly using the CoT format inside the exponential function in Eq.4 by detecting it from motion reward fails to generate stable policies via end-to-end training. As a result, authors in [5] learn a contact schedule instead and execute it via low-level MPC. In our approach, we deployed the exponential form, which guarantees a positive reward and scale it within $(0,1]$. We multiply the absolute values of each entry of $\tau$ with each entry of $\dot{q}$ and sum them up in (4) to follow the fact that a motor does not get charged back even when the applied torque is opposite to the motion [33]. This adaptive energy regularization term allows us to learn stable locomotion policies across different speeds, tasks, and embodiments.
One may claim that the selection of an appropriate energy reward weight $\alpha_{en}$ in (2) is still nontrivial since a tiny $\alpha_{en}$ diminishes the influence of energy regularization. At the same time, an overly large value can lead to over-enforcement, compromising velocity tracking accuracy. However, the generalized distance design calibrates the energy reward across different speeds, and the exponential function design on motion and energy reward scale both into $(0,1]$. As a result, a $\alpha_{en}=1$ works for all cases in our experiments, and the parameter is not that sensitive. This trade-off and detailed ablation study will be shown in Section IV.
III-3 Auxiliary Rewards
Energy regularization alone is usually insufficient for generating proper behaviors. To address this, we add a few auxiliary regularization rewards, including collision avoidance, action rate control and trunk orientation regularization. Higher values of these terms signify less desirable performance, so we deployed a negative exponential function as detailed in (2). Comprehensive information on these auxiliary rewards can be found in the project website.
IV Experiments
The experiments are designed to support the following statements after adding the velocity-dependent energy reward as shown in (4) without specifying gait information.
  • The generated RL policy automatically selects an energy efficient action at different command linear and angular velocity.
  • The energy reward weight $\alpha_{en}$ should be comparable to motion rewards in order to get a satisfactory RL policy.
  • The energy reward can also be used for locomotion training on terrains with minor amendments on the auxiliary reward $R_{aux}$.
IV-A Training Setup